OpenStudy (anonymous):

HELP!!! In a 45°-45°- 90° triangle, if the legs each have a length of 4, then what is the length of the hypotenuse? A. the square root of2 B. 4the square root of2 C. 4 D. 4the square root of3

OpenStudy (darkigloo):

the 45-45-90 triangle is one of those special right triangles. in this triangle this rule applies: hypotenuse = (leg)(√2) If you remember that rule, you can easily find the hypotenuse. hypotenuse = 4 √2 You can also use the Pythagorean Theorem. a^2 + b^2 = c^2 and you'll get the same answer.

OpenStudy (anonymous):

so its b?

OpenStudy (darkigloo):

yeah